OpenStudy (anonymous):

There is a total of $620 in tips collected over an 8-hour period for 8 employees. Four worked 8 hours; four worked 5 hours. What does each get in tips so it equals $620

OpenStudy (bobbyleary):

This is actually applicable to my job and this is how we do it. First find the total amount of hours worked for the whole day. There were four that worked 8 hours: 4(8) = 32 There were four that worked 5 hours: 4(5) = 20 To find the total amount of hours worked, you add these totals. 32+20 = 52 Now to find an hourly rate ($/hr), take the total amount of tips divided by the total amount of worked hours. Rate = $620/52hr = $11.92/hr Now to find out how much each employee earned, you multiply this rate by the number of hours THEY worked. For the people who worked 8 hours: (8hr)*($11.92/hr) = $95.36 For the people who worked 5 hours: (5hr)*($11.92/hr) = $59.60

OpenStudy (bobbyleary):

And to double check, 4*($95.36) + 4*($59.60) = $619.84 This is slightly off due to rounding the hourly rate. For a more exact result, use one or two more decimal places for the hourly rate, such as $11.9230.

OpenStudy (anonymous):

Thank you, Thank you, Thank you--you have now idea how many of us could not get this to come out.

OpenStudy (bobbyleary):

No problem. I'm glad I could help!